Recently, More Money 4U offered an annuity that pays 6.3% compounded monthly. If $1,848 is deposited into this annuity every​ month, how much is in the account after ​years? How much of this is​ interest?
1
Expert's answer
2020-11-24T11:20:50-0500

Let the term be 3 years, than we solved by formula:

"FV=A\times\frac{(1+i)^n-1}{i}"


"i=\frac{0.063}{12}=0.00525"


"n=3*12=36"


"FV=A\times\frac{(1+i)^n-1}{i}=1848\times\frac{(1+0.00525)^{36}-1}{0.00525}=73020.24"


interest

"73020.24-1848\times36=73020.24-66528=6492.24"